# Towards Distributed Backdoor Attacks with Network Detection in Decentralized Federated Learning

### v1

@@ Abstract
With the rapid development of deep learning models, increasing attention has been paid to concerns about user privacy and copyright Hallaji et al. (2024); Xiao et al. (2025); Nguyen et al. (2022). Federated learning (FL) (McMahan et al., 2017; Kairouz et al., 2021; Bai et al., 2024) is a promising paradigm for collaborative training machine learning models over large-scale distributed data. Based on the communication architecture, existing FL frameworks can be classified into two categories: centralized FL and decentralized FL (Li et al., 2023b). Specifically, in centralized FL, the server updates the global model by aggregating the information from parties (McMahan et al., 2017; Li et al., 2020; Wang et al., 2024; Hamer et al., 2020). In decentralized FL, the communications are performed among the parties and every party can update the global parameters directly (Bornstein et al., 2023; Marfoq et al., 2020; Shi et al., 2023; Dai et al., 2022)
@@
Figure 2: Attacks on D-PSGD
In this paper, we aim to achieve a high attack success rate regardless of the locations of adversarial agents. First, we propose to detect the network by predicting the distance between any two attackers on the network. Specifically, we observe that the sequence of prediction accuracy of elaborated data varies differently on agents with different distances to an attacker. Based on this observation, we use the sequence to predict the distance between any two attackers in the early stage of FL. With the estimated distance, we leverage the clustering algorithm to organize the attackers in different clusters. Lastly, we develop an algorithm to dynamically decompose global trigger patterns into different adversarial agents to maximize the attack success rate. Compared with DBA, our method has addressed the distinctive framework of decentralized FL and achieved a higher attack success rate.
We experiment with multiple decentralized FL frameworks and standard datasets to verify the effectiveness of the proposed method. In summary, we propose the following contributions:
  • This work is the first to study distributed backdoor attacks on decentralized FL. We empirically find that while directly applying DBA to decentralized FL, the attack success rate depends on the distribution of attackers in the topology of decentralized FL.
  • We propose a method to detect the network of the decentralized FL by estimating the distance between any two agents. An algorithm is developed to dynamically organize distributed backdoor attacks based on clusters.
@@ 1 Preliminary
1.1 Federated Learning
Centralized Federated Learning (FL) is a distributed learning framework with the following training objective:
$\vspace{-1mm}\min_{w}F(w):=\frac{1}{N}\sum^{N}_{i=1}f_{i}(w_{i})$ (1)
There are $N$ parties in the framework, each of whom trains a local model $f_{i}(w)$ with a private dataset $D_{i}=\{\{x_{j}^{i},y_{j}^{i}\}_{j=1}^{J}\}$ where $j=|D_{i}|$ and $\{x_{j}^{i},y_{j}^{i}\}$ represents each data sample and its corresponding label. At round t, a central server sends the current shared model parameterized with $w$ to $N$ parties. Each local party will copy $w$ to its local model $w_{i}$. The parameter of a local model $w_{i}$ will be updated with a loss of prediction $l(\{\{x_{j}^{i},y_{j}^{i}\}_{j=1}^{J}\},w_{i})$. By running an optimization algorithm such as stochastic gradient descent, a local party can obtain a new local model $w_{i}^{t+1}$. After several rounds, the server implements an aggregation algorithm to combine the local models or model updates into a global model which is then disseminated back to the local parties:
$w^{t+1}=w^{t}+\frac{\eta}{N}\sum^{N}_{i=1}(w_{i}^{t+1}-w^{t}),$ (2)
where $\eta$ is the parameter to decide the step size of the update. This distributed learning framework preserves data privacy by training models locally on distributed devices. Instead of sharing actual data with a central server, only local models or local model updates are shared. The averaging algorithm can also be replaced by other algorithms such as FedMedian (Yin et al., 2018).
@@ 1.1 Federated Learning
1.2 Backdoor attack
The objective of a backdoor attack is to mislead the trained model to predict any input data with an embedded trigger as a wrong label. In federated learning, an adversarial client can pretend to be a normal client and manipulate the local model. By sending the updates to the global server or neighbors, the global model would achieve a high attack success rate on poisoned data while behaving normally on clear data samples to fit the main task. Specifically, the training objective for an adversarial client $i$ at round $t$ with local dataset $D_{i}$ and the target label $\tau$ is:
$w_{i}^{*}=\operatorname*{arg\,max}_{w_{i}}(\sum_{j\in S_{\text{poi}}^{i}}P[F(w,R(x_{j}^{i}))=\tau]+\sum_{j\in S_{\text{cln}}^{i}}P[F(w,x_{j}^{i})=y_{j}^{i}]),$ (3)
@@ 2 Method
2.1 Analysis of DBA in Decentralized Federated Learning
@@ 2.2 Topology Detection
Figure 3: Sequences
To verify that the signal is useful for distance prediction, we first start with a simple experiment: given a sequence of poison accuracy in the training process, we use a Long Short-Term Memory network (LSTM) model to predict if it comes from the attacker or not (binary classification). On a decentralized FL configured with a ring topology with 8 nodes, where a client performs the attack. The training process generates 8 sequences for all 8 nodes. We find that the experimental accuracy reached 100%. The sequence from the attacker exhibits significant temporal differences from other clients.
@@ 3 DBA based on the detected network
$w_{i}^{*}=\operatorname*{arg\,max}_{w_{i}}(\sum_{j\in S_{\text{poi}}^{i}}P[F(w,R(x_{j}^{i},\phi_{k}^{i}))=\tau]+\sum_{j\in S_{\text{cln}}^{i}}P[F(w,x_{j}^{i})=y_{j}^{i}]),$ (6)
@@
Similar to DBA, there are multiple factors to be explored in decentralized FL: location, size, and gap. The location is the offset of the trigger pattern from the top left pixel. The Size decides the number of pixels of the trigger covered. Trigger Gap is used to shift the local trigger from the previous local trigger. Instead of continuously attacking throughout the training process, we considered allowing the attacker to attack at intervals. This is similar to alternating current where the signal strength transmitted to each node will vary over time, and periodic decreases in signal strength could make the fluctuations more apparent.
Algorithm 1 outlines the workflow of our attacking scheme. In the early stage of learning ($t<\Delta T$), the sequence of poison accuracy will be used for predicting the distance between any two attackers. Based on the distance matrix, we can leverage any clustering algorithm based on distance to group attackers. Then in each attacking round, the global trigger will be decomposed and randomly assigned to all attackers within each cluster. Each attacker will conduct backdoor attacks with the assigned local trigger. Our cluster-based on backdoor attacks and dynamic distribution of local triggers can enhance the impact of distributed backdoor attacks.
$t=0$;
Assign a distinctive poison signature out of the domain for each attacker;
while $t<\Delta T$ do
       for $i\in\mathcal{A}$ do
             for $i\in\mathcal{A}$ do
                   Compute the poison accuracy $s_{i}$ for attacker $i^{\prime}$ concerning $z_{i}$;
             end for
       end for
      t+=1;
end while
For any pair of two attackers, predict the distance $d_{i,i^{\prime}}$ from $i$ to $i^{\prime}$ with $G(\cdot)$, $s_{i}$, and $s_{i^{\prime}}$;
Clustering attackers into $K$ groups with the distance matrix $M$;
while $t<T$ do
       for $k=0;\ k<K;\ k+=1$ do
             Randomly assign decomposed local patterns to all attackers $i$ in Cluster $C_{k}$;
             for $i\in C_{k}$ do
                   Each attacker $i$ uses Eq. (6) to attack the local model;
             end for
       end for
      t+=1;
end while
Algorithm 1 DBA with network detection
@@ 4 Experiments
Experimental Setup We follow DBA Xie et al. (2020) to set up the experiment. We introduce two popular decentralized FL algorithms: DSGD Amiri & Gündüz (2020) and Swift Bornstein et al. (2023). All training parameters are configured as the standard value in the corresponding paper. We evaluate the performance of predicting distance on two typologies: Ring and Grid. To compare with DBA and centralized backdoor attack Bagdasaryan et al. (2020), we report the attack success rate (ASR) on two datasets: CIFAR-10 and MNIST. We use the poison accuracy of the first 100 epochs to predict distance. On each topology, there are 40 clients by default. We follow DBA to set up the attacking trigger.
@@
(c) Grid Topology (DSGD)
(d) Grid Topology (Swift)
@@
Distance prediction. In the experiment, we randomly assign pairs of clients as attackers with specified ground-truth distance and leverage an LSTM model to predict distance. The experiments are repeated 20 times to combat randomness. As shown in Figure 5, we report the error of the predicted distance on two typologies with different numbers of clients. On the ring topology, we can observe the prediction error for Swift is smaller than DSGD. We attribute it to the rapid synchronization of model updates in Swift. The observations still hold for the grid topology. Also, we can see that the error increases while the ground-truth distance is increasing. This is because the attacking signal becomes weak if the distance is long and the model can not distinguish it from the signal of a non-attack client. The overall result indicates that our distance prediction method is accurate.
The robustness of distributed backdoor attack. Following DBA, we evaluate the attack success rates of different attacking methods using the same global trigger. The ratio of backdoor pixels in the global triggers is 0.964 for MNIST and 0.990 for CIFAR-10. For a fair comparison, we set the total number of backdoor pixels in the training dataset to be the same across different attacking methods. Specifically, we poison more data in DBA and centralized attack so that the total number of poison pixels equals that of our cluster-based DBA by including more data in $S^{i}_{poi}$. We randomly select 10 clients as attackers and cluster the attackers into 3 groups. In Figure 7, we use “Cluster-based DBA” to denote our method. We report the average attack success rate for two topologies on CIFAR-10 and MNIST. We can see that the centralized attack outperforms DBA in terms of attack success rate. This is against the motivation of distributed backdoor attacks in FL. It further justifies the necessity of improving DBA in decentralized FL.
@@
Figure 6: Ring Topology
@@
Figure 7: Clique Ring Topology
@@
By applying our backdoor attack method to the decentralized FL, we can observe that our attack success rate is higher than both DBA and centralized attacks in all settings. Specifically, on the CIFAR-10 dataset, the attack success rate is always higher than DBA and centralized attacks in terms of both two topologies with two decentralized FL frameworks. On the MNISIT dataset, we observe that the success attack ratios of all three methods reach 100% in the end. However, our method has a higher ASR in the search stage and converges faster than the other two methods. The superiority of our method over DBA and centralized attacks in various settings verifies that our strategy can address the limitation of DBA in decentralized FL.
@@
Case study. In Figure 8, we use the Grad-CAM visualization method Gildenblat & contributors (2021) to explore a sample image attacked by DBA and centralized attack with the global trigger. The two columns show the difference between two heat maps of activation (e.g., the importance for prediction) for predicting a hand-written digit ‘4’ as ‘4’ and ‘2’, respectively. Same as the conclusion in DBA Xie et al. (2020), each local triggered image alone is a weak attack as none of them can change the prediction. However, with a global trigger, the poisoned image is classified as ‘2’ (the target label), and we can see the activation area is transformed to the trigger location. It suggests that each small local trigger is difficult to detect for defenders because most locally triggered images are similar to the clean image, demonstrating the stealthy nature of distributed backdoor attacks. We remark that distributed attacks can make triggers stealthier. In the following table, we use the strategy in Zhang et al. (2022) to attack parameters in DFL. The results in Figure 1 of Appendix indicate that DBA can further increase durability. This is because each decomposed trigger is small and it makes the one-line gradient project in [1] more invisible. We totally agree that finding an optimal combination of many parameters is a challenge for DBA. Our contribution is that for any combination of the parameters, our clustering algorithm can improve the attack success rate.
We also follow DBA to investigate the effects of trigger factors in the process of decomposing a global trigger. We only change one factor in each experiment shown in Figure 9. When we increase the size of the local trigger from 1 to 4, the attack success ratio will increase. At the same time, the accuracy varies slightly. However, while increasing the size from 4 to 12, the attack success ratio will drop. The value of the gap has little impact on both ASR and accuracy. This is because the relation between different local triggers has been removed by distributing the local triggers to different clients. We also note a U-shape curve of ASR when the shift increases. This is because when the trigger overlaps with some pattern in the clear image, the impact can be ignored due to overlap. However, when we further shift the trigger to the right bottom corner, the ASR will recover to a high ratio because most objects are located in the middle of the images in the dataset.
@@
5 Related works
Using more data for model training benefit the performance in general. However, it poses privacy risk concerns by collecting data from various institutions. Federated Learning McMahan et al. (2017); Khaled & Jin (2023); Cheng et al. (2024); Zhong et al. (2023) has emerged as a powerful distributed learning framework by sharing a global model without sharing their data. FL frameworks can be classified into two categories: centralized FL and decentralized FL Li et al. (2023b). Centralized FL enables clients to perform limited training on local datasets while the centralized server aggregates the client parameters using different aggregation methods. (McMahan et al., 2017; Li et al., 2020; Wang et al., 2024; Hamer et al., 2020; Murad et al., 2024). In decentralized FL, the communications are performed among the parties and every party can update the global parameters directly (Bornstein et al., 2023; Marfoq et al., 2020; Shi et al., 2023; Dai et al., 2022) to keep each client’s data private.
The nature of federated learning provides a way for adversarial parties to attack the model because any client on the communication topology can pretend to be a normal client and manipulate the model update by injecting poison data. Since any client has access to the global model, the attacker can perform membership attacks on the model (Li et al., 2023a), data stealing (Garov et al., 2024) or model poisoning attack Yan et al. (2023); Jia et al. (2023); Zhang et al. (2023a); Li et al. (2022); Huang et al. (2021). Some defensive methods have also been studies (Xie et al., 2024; 2021; Zhang et al., 2023b) based on model updates. However, the attack and defense on the decentralized FL have not been studied. To the best of our knowledge, our paper is the first work to investigate distributed backdoor attacks on decentralized FL.
6 Conclusion
In this paper, we apply DBA to decentralized FL. We experimentally demonstrate that the attack success rate of DBA depends on the distribution of attackers in the network architecture. Considering that the attackers can not decide their location, we propose a two-step attacking strategy to improve the ASR of DBA in decentralized FL: (1) detecting the network and (2) an improved DBA based on the network. Lastly, we propose an algorithm to dynamically embed local patterns decomposed from a global pattern into the different attackers in each cluster. We experimentally verify that our attacking strategy can achieve a higher attack success rate than DBA and the centralized attack.
Appendix A Appendix
Table 1:
Method Swift-Ring DSGD-Ring Swift-Clique DSGD-Clique
Neurotoxin 0.601 0.623 0.652 0.672
Neurotoxin+DBA 0.613 0.636 0.662 0.663
Neurotoxin+our 0.651 0.676 0.712 0.704
A.1 Attacking DFL with defensive mechanisms
To showcase the effectiveness of the proposed attack under defense mechanisms, we introduce two defensive mechanisms Zhang et al. (2023b); Jia et al. (2023) in the experiment. To the author’s best knowledge, there is no defense mechanism designed specifically for decentralized FL in the literature. The possible reason is that a decentralized framework itself is a defense mechanism. Many defensive strategies based on client selection such as Krum Blanchard et al. (2017) are not suitable for DFL. To introduce the defensive strategy in decentralized FL, we leverage the corresponding strategy for each client. Note that the defense mechanism does reduce ASR. However, decentralized FL mitigates backdoor attacks because each client only has a few neighbors (e.g., 2 on a ring topology). Compared with DBA and centralized attack, our method can further pose a challenge to the effectiveness of these defense mechanisms.
Table 2: Attacking DFL with defensive mechanism
Method DBA Centralized Our
Swift 0.656 0.782 0.801
Swift+FLIP 0.431 0.699 0.783
Swift+FedGame 0.587 0.728 0.779
DSGD 0.712 0.764 0.831
DSGD+FLIP 0.679 0.688 0.787
DSGD+FedGame 0.646 0.647 0.805

### v2

# DBA-DFL: Towards Distributed Backdoor Attacks with Network Detection in Decentralized Federated Learning

@@ Abstract
Machine Learning, ICML
1 Introduction
Federated learning (FL) (McMahan et al., 2017; Kairouz et al., 2021; Bai et al., 2024) is a promising paradigm for collaborative training machine learning models over large-scale distributed data. It preserves the privacy of local data in each client and enjoys the advantage of efficient optimization as the local clients conduct computations independently and simultaneously (Andrew et al., 2024). Based on the communication architecture, existing FL frameworks can be classified into two categories: centralized FL and decentralized FL (Li et al., 2023b). Specifically, in centralized FL, the server updates the global model by aggregating the information from parties (McMahan et al., 2017; Li et al., 2020b; Wang et al., 2024; Hamer et al., 2020). In decentralized FL, the communications are performed among the parties and every party can update the global parameters directly (Bornstein et al., 2023; Li et al., 2020a; Marfoq et al., 2020; Shi et al., 2023; Dai et al., 2022)
@@ 1 Introduction
Figure 1: Location of Attackers
Figure 2: Attacks on D-PSGD
@@
In this paper, we aim to achieve a high attack success rate regardless of the locations of adversarial agents. First, we propose to detect the network by predicting the distance between any two attackers on the network. Specifically, we observe that the sequence of prediction accuracy of elaborated data varies differently on agents with different distances to an attacker. Based on this observation, we use the sequence to predict the distance between any two attackers in the early stage of FL. With the estimated distance, we leverage the clustering algorithm to organize the attackers in different clusters. Lastly, we develop an algorithm to dynamically decompose global trigger patterns into different adversarial agents to maximize the attack success rate. Our method has addressed the distinctive framework of decentralized FL and achieved a higher attack success rate.
We experiment with multiple decentralized FL frameworks and standard datasets to verify the effectiveness of the proposed method. We propose the following contributions:
  • This work is the first to study distributed backdoor attacks on decentralized FL.
  • We empirically find that while directly applying DBA to decentralized FL, the attack success rate depends on the distribution of attackers in the topology.
  • We propose a method to detect the network of the decentralized FL by estimating the distance between any two agents. An algorithm is developed to dynamically organize distributed backdoor attacks.
@@ 2 Preliminary
Federated Learning. Centralized FL is a distributed learning framework with the following training objective:
$\vspace{-2mm}\min_{w}F(w):=\frac{1}{N}\sum^{N}_{i=1}f_{i}(w_{i})$ (1)
There are $N$ parties in the framework, each of whom trains a local model $f_{i}(w)$ with a private dataset $D_{i}=\{\{x_{j}^{i},y_{j}^{i}\}_{j=1}^{J}\}$ where $j=|D_{i}|$ and $\{x_{j}^{i},y_{j}^{i}\}$ represents each data sample and its corresponding label. At round t, a central server sends the current shared model parameterized with $w$ to $N$ parties. Each local party will copy $w$ to its local model $w_{i}$. The parameter of a local model $w_{i}$ will be updated with a loss of prediction $l(\{\{x_{j}^{i},y_{j}^{i}\}_{j=1}^{J}\},w_{i})$. By running an optimization algorithm such as stochastic gradient descent, a local party can obtain a new local model $w_{i}^{t+1}$. After several rounds, the server implements an aggregation algorithm to combine the local models or model updates into a global model.
@@
Backdoor attack The objective of a backdoor attack is to mislead the trained model to predict any input data with an embedded trigger as a wrong label. In federated learning, an adversarial client can pretend to be a normal client and manipulate the local model. By sending the updates to the global server, the global model would achieve a high attack success rate on poisoned data. Specifically, the training objective for an adversarial client $i$ at round $t$ with local dataset $D_{i}$ and the target label $\tau$ is:
$\begin{split}w_{i}^{*}=&\operatorname*{arg\,max}_{w_{i}}(\sum_{j\in S_{\text{poi}}^{i}}P[F(w,R(x_{j}^{i}))=\tau]\\ &+\sum_{j\in S_{\text{cln}}^{i}}P[F(w,x_{j}^{i})=y_{j}^{i}]),\end{split}$ (2)
@@ 3 Method
3.1 Analysis of DBA in Decentralized FL
@@ 4 DBA based on the detected network
$\begin{split}w_{i}^{*}=&\operatorname*{arg\,max}_{w_{i}}(\sum_{j\in S_{\text{poi}}^{i}}P[F(w,R(x_{j}^{i},\phi_{k}^{i}))=\tau]\\ &+\sum_{j\in S_{\text{cln}}^{i}}P[F(w,x_{j}^{i})=y_{j}^{i}]),\end{split}$ (5)
@@
1 $t=0$;
2 Assign a distinctive poison signature out of the domain for each attacker;
3 while $t<\Delta T$ do
4       for $i\in\mathcal{A}$ do
5             for $i\in\mathcal{A}$ do
6                   Compute the poison accuracy $s_{i}$ for attacker $i^{\prime}$ concerning $z_{i}$;
7
8      t+=1;
9For any pair of two attackers, predict the distance $d_{i,i^{\prime}}$ from $i$ to $i^{\prime}$ with $G(\cdot)$, $s_{i}$, and $s_{i^{\prime}}$;
10 Clustering attackers into $K$ groups with the distance matrix $M$;
11
12while $t<T$ do
13       for $k=0;\ k<K;\ k+=1$ do
14             Randomly assign decomposed local patterns to all attackers $i$ in Cluster $C_{k}$;
15             for $i\in C_{k}$ do
16                   Each attacker $i$ uses Eq. (5) to attack the local model;
17
18
19      t+=1;
Algorithm 1 DBA with network detection
@@ 14             Randomly assign decomposed local patterns to all attackers $i$ in Cluster $C_{k}$;
Algorithm 1 outlines the workflow of our attacking scheme. In the early stage of learning ($t<\Delta T$), the sequence of poison accuracy will be used for predicting the distance between any two attackers. Based on the distance matrix, we can leverage any clustering algorithm based on distance to group attackers. Then in each attacking round, the global trigger will be decomposed and randomly assigned to all attackers within each cluster. Each attacker will conduct backdoor attacks with the assigned local trigger. Our cluster-based on backdoor attacks and dynamic distribution of local triggers can enhance the impact of DBA.
@@ 5 Experiments
Experimental Setup We follow DBA (Xie et al., 2020) to set up the experiment. We introduce two popular decentralized FL algorithms: DSGD (Amiri & Gündüz, 2020) and Swift (Bornstein et al., 2023). All training parameters are configured as the standard value in the corresponding paper. We evaluate the performance of predicting distance on two typologies: Ring and Grid. To compare with DBA and centralized backdoor attack (Bagdasaryan et al., 2020), we report the attack success rate (ASR) on two datasets: CIFAR-10 and MNIST. We use the poison accuracy of the first 100 epochs to predict distance. On each topology, there are 40 clients by default. We follow DBA to set up the attacking trigger.
@@
(c) Grid Topology (Swift)
@@
Figure 6: Attack Success Rate (Ring Topology)
@@
Figure 7: Attack Success Rate (Clique Ring Topology)
Distance prediction. In the experiment, we randomly assign pairs of clients as attackers with specified ground-truth distance and leverage an LSTM model to predict distance. The experiments are repeated 20 times to combat randomness. As shown in Figure 5, we report the error of the predicted distance on two typologies with different numbers of clients. On the ring topology, we can observe the prediction error for Swift is smaller than DSGD. We attribute it to the rapid synchronization of model updates in Swift. The observations still hold for the grid topology. Also, we can see that the error increases while the ground-truth distance is increasing. This is because the attacking signal becomes weak if the distance is long and the model can not distinguish it from the signal of a non-attack client. The result indicates that our distance prediction method is accurate.
Attack success rate. Following DBA, we evaluate the attack success rates of different attacking methods using the same global trigger. The ratio of backdoor pixels in the global triggers is 0.964 for MNIST and 0.990 for CIFAR-10. For a fair comparison, we set the total number of backdoor pixels in the training dataset to be the same across different attacking methods. Specifically, we poison more data in DBA and centralized attack so that the total number of poison pixels equals that of our cluster-based DBA by including more data in $S^{i}_{poi}$. We randomly select 10 clients as attackers and cluster the attackers into 3 groups. In Figure 7, we use “Cluster-based DBA” to denote our method. We report the average attack success rate for two topologies on CIFAR-10 and MNIST. We can see that the centralized attack outperforms DBA in terms of attack success rate. This is against the motivation of DBA. It further justifies the necessity of improving DBA in decentralized FL. By applying our backdoor attack method to the decentralized FL, we can observe that our attack success rate is higher than both DBA and centralized attacks in all settings.
Attacking DFL with defensive mechanisms. To showcase the effectiveness of the proposed attack under defense mechanisms, we introduce two defensive mechanisms (Zhang et al., 2023b; Jia et al., 2023) in Table 1. To the author’s best knowledge, there is no defense mechanism designed specifically for decentralized FL in the literature. The possible reason is that a decentralized framework itself is a defense mechanism. Many defensive strategies based on client selection such as Krum (Blanchard et al., 2017) are not suitable for DFL. To introduce the defensive strategy in decentralized FL, we leverage the corresponding strategy for each client. Note that the defense mechanism does reduce ASR. However, decentralized FL mitigates backdoor attacks because each client only has a few neighbors (e.g., 2 on a ring topology). Compared with DBA and centralized attack, our method can further pose a challenge to the effectiveness of these defense mechanisms.
Table 1: Attacking DFL with defensive mechanism.
Method DBA Centralized Ours
Swift 0.656 0.782 0.801
Swift+FLIP 0.431 0.699 0.783
Swift+FedGame 0.587 0.728 0.779
DSGD 0.712 0.764 0.831
DSGD+FLIP 0.679 0.688 0.787
DSGD+FedGame 0.646 0.647 0.805
@@
Table 2: Number of clusters.
Clusters Swift-Ring DSGD-Ring Swift-Clique
2 clusters 0.789 0.812 0.872
3 clusters 0.801 0.831 0.893
4 clusters 0.818 0.823 0.876
5 clusters 0.752 0.788 0.862
@@
Case study. In Figure 8, we use the Grad-CAM visualization method (Gildenblat & contributors, 2021) to explore a sample image attacked by DBA and centralized attack with the global trigger. The two columns show the difference between two heat maps of activation (e.g., the importance for prediction) for predicting a hand-written digit ‘4’ as ‘4’ and ‘2’, respectively. Same as the conclusion in DBA (Xie et al., 2020), each local triggered image alone is a weak attack as none of them can change the prediction. However, with a global trigger, the poisoned image is classified as ‘2’ (the target label), and we can see the activation area is transformed to the trigger location. It suggests that each small local trigger is difficult to detect for defenders because most locally triggered images are similar to the clean image.
6 Related works
Using more data for model training benefit the performance in general. However, it poses privacy risk concerns by collecting data from various institutions. Federated Learning (McMahan et al., 2017; Khaled & Jin, 2023; Cheng et al., 2024; Huang et al., 2021a; Zhong et al., 2023) has emerged as a powerful distributed learning framework by sharing a global model without sharing their data. FL frameworks can be classified into two categories: centralized FL and decentralized FL (Li et al., 2023b). Centralized FL enables clients to perform limited training on local datasets while the centralized server aggregates the client parameters using different aggregation methods. (McMahan et al., 2017; Li et al., 2020b; Wang et al., 2024; Hamer et al., 2020). In decentralized FL, the communications are performed among the parties and every party can update the global parameters directly (Bornstein et al., 2023; Li et al., 2020a; Marfoq et al., 2020; Shi et al., 2023; Dai et al., 2022) to keep each client’s data private.
The nature of federated learning provides a way for adversarial parties to attack the model. Since any client has access to the global model, the attacker can perform membership attacks on the model (Li et al., 2023a), data stealing (Garov et al., 2024) or model poisoning attack (Yan et al., 2023; Jia et al., 2023; Zhang et al., 2023a; Li et al., 2022; Huang et al., 2021b). Some defensive methods have also been studies (Xie et al., 2024, 2021; Zhang et al., 2023b; Fang & Chen, 2023) based on model updates. However, the attack and defense on the decentralized FL have not been studied. To the best of our knowledge, our paper is the first work to investigate DBA on decentralized FL.
7 Conclusion
In this paper, we apply DBA to decentralized FL. We experimentally demonstrate that the attack success rate of DBA depends on the distribution of attackers in the network architecture. Considering that the attackers can not decide their location, we propose a two-step attacking strategy to improve the ASR of DBA in decentralized FL: (1) detecting the network and (2) an improved DBA.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, such as safety and privacy in federated learning.
Appendix A Appendix
@@
Discussion of parameters. We conduct experiments to vary the number of clusters. We use K-means as the clustering algorithm. So the cluster size will automatically decided by the value of K. The results in Table 2 suggest that there is a tradeoff in choosing the value of K. When K is too small, it tends to be similar to DBA. When K is too large, it is similar to centralized attacks. We investigate the impact of the error of distance prediction on the method’s effectiveness in Figure 10. Since we can directly control inaccuracies in distance prediction, we vary the topology of DFL and the number of clients. With more clients and random structures, we have observed larger errors. The following table shows ASR when the error is varying. We have not observed an error larger than 5.3. Even in the worst case, ASR with our method is still higher than DBA. We remark that it is unnecessary to set a threshold because our prediction will never be worse than random distribution in DBA. We also follow DBA to investigate the effects of trigger factors in the process of decomposing a global trigger. We only change one factor in each experiment shown in Figure 9. When we increase the size of the local trigger from 1 to 4, the attack success ratio will increase. At the same time, the accuracy varies slightly. However, while increasing the size from 4 to 12, the attack success ratio will drop. The value of the gap has little impact on both ASR and accuracy. This is because the relation between different local triggers has been removed by distributing the local triggers to different clients. We also note a U-shape curve of ASR when the shift increases. This is because when the trigger overlaps with some pattern in the clear image, the impact can be ignored due to overlap. However, when we further shift the trigger to the right bottom corner, the ASR will recover to a high ratio because most objects are located in the middle of the images in the dataset.
Table 3: Computational cost of our method.
Topology Extra cost 2000 epochs of training
40 nodes 9 minutes 3 hours
80 nodes 25 minutes 9 hours
100 nodes 32 minutes 11 hours
Figure 10: The Effects of Distance Prediction Error
Computational cost. In table 3, we report the running time of our algorithm and the training time of FL. We remark that the majority of the computational overhead is still the cost of training on the decentralized FL. For Cifar-100, it usually takes at least 3000 epochs to reach the convergent performance with FL. The cost of clustering can be ignored compared with training. Also, trigger distribution can be done in a few seconds. Therefore, our method can handle more complex real-world topologies, and the extra computational overhead can be ignored.